OpenStudy (anonymous):

a 2 kg rubber ball moving at 6m/s to the right hits a stationary 6 kg. after the collision, the 2 kg ball has a velocity of 3m/sec in the opposite direction. find the velocity of the 6kg ball

OpenStudy (anonymous):

momentom before and after the collision have to be equal. so you get m1v1+m2v2 = momentum total (2)(6)+(6)(0)=12 knowing that it has to equal 12 and knowing your new speed of the 2kg ball you get (2)(3)+(6)(x)=12 x=(12-6)/6 x=1

OpenStudy (anonymous):

sorry i need to make a change to that give me a min

OpenStudy (anonymous):

my answer sheet says 3 @roblsander

OpenStudy (anonymous):

(2)(-3)+(6)(x)=12 x=(12+6)/6 x=3

OpenStudy (anonymous):

i forgot to include that the 2kg was moving opposite compared to before which makes it -3m/s instead of 3m/s

OpenStudy (anonymous):

thanks @roblsander

OpenStudy (anonymous):

No problem

OpenStudy (anonymous):

can you help me with another problem? @roblsander

OpenStudy (anonymous):

depends what it is ill try

OpenStudy (anonymous):

two rubber balls collide in a collision, one ball has a mass of 2 kg and has a velocity of 6m/s due east. the other ball has a mass of 6 kg and is traveling behind the other ball at a velocity of 8 m/sec due east. after the collision, the 6 kg ball slows down to 7 m/sec . find the velocity of the 2 kg ball after the collision

OpenStudy (anonymous):

@roblsander

OpenStudy (anonymous):

\[P_t=m_1v_1+m_2v_2\] \[P_b=(2)(6)+(6)(8)\] \[P_b=60\] \[P_b=P_a\] \[60=(6\times7)+(2x) \] \[x=9\]